# When ControlNet Meets Inexplicit Masks: A Case Study of ControlNet on its Contour-following Ability

### v2

@@ 4.2 Impact of Inexplicit Masks
To clarify concerns about ControlNet’s performance on inexplicit masks, we first imitate inaccurate control signals by dilating object masks $\boldsymbol{m}$ with progressive radius $r(r\geq 0)$. The dilated mask is denoted as $\boldsymbol{m}_{r}$, where $\boldsymbol{m}_{0}$ presents masks with precise contours. Considering that bounding-box masks only convey object position and size without identifiable contours, we take an intersection over the dilated mask $\boldsymbol{m}_{r}$ and the bounding-box mask, denoted as $\boldsymbol{b}$, to construct inexplicit masks of various precision, formulated as,
$\boldsymbol{m}_{r}=Dilate(\boldsymbol{m},r)\cap\boldsymbol{b},$ (4)
@@
where the bounding-box mask $\boldsymbol{b}$ can also be expressed as the extreme case $\boldsymbol{m}_{\infty}$ ($\boldsymbol{b}\subset\boldsymbol{m}_{\infty}$). In the rest of the paper, we use $\boldsymbol{m}_{\infty}$ to denote the bounding-box mask.
@@
We use the notation ControlNet-$\boldsymbol{m_{r}}$ to denote ControlNet trained with deteriorated masks $\boldsymbol{m}_{r}$, where the vanilla ControlNet trained with precise masks is ControlNet-$\boldsymbol{m_{0}}$. We test these models on masks at varying deterioration degrees to explore the performance of ControlNet with inexplicit masks and its contour-following ability. Fig. 2 presents the experimental results. Here are two main observations.
@@
The vanilla ControlNet, i.e., ControlNet-$\boldsymbol{m}_{0}$, strictly adheres to the provided outlines, where inexplicit masks would cause severe performance degradation. As shown in Fig. 1, the generated images are faithful to the contours of the control inputs under various control modalities, which is the key to excellent spatial controllable generation. One underlying reason is supposed to be that edges provide a shortcut to reconstruct input images. Therefore, ControlNet exhibits strong contour instructions during the generation process.
@@
However, such a contour-following ability takes no account for inaccurate contours and maintains a high average CR of over 60% in Fig. 2. While it allows minor editing by manipulating masks to some extent, this property imposes severe problems on image realism when inaccurate contours are provided. Similar observations are also noted in Voynov et al. (2023); Bhat et al. (2023); Singh et al. (2023). For example, ControlNet tends to interpret rectangle masks as boards rather than animals specified in the prompt, causing incorrect spatial layouts and distorted objects. It is also reflected in a degradation of 1.35 CLIP-score and 8.62 FID. Thus, precise masks with accurate contours are necessary for high-quality generation. However, obtaining masks with accurate contours is challenging, especially for non-expert users.
@@
Training ControlNet with deteriorated masks also converges and exhibits higher robustness on masks of varying precision with a sacrifice of contour instructions. We further train a series of ControlNet on manually deteriorated masks. Surprisingly, we find ControlNet converges successfully with all conditions, even on the extreme bounding-box masks. As illustrated in Fig. 2, training with deteriorated masks consistently improves ControlNet’s robustness in interpreting inexplicit masks with better CLIP-score and FID at the cost of weak contour-following ability.
@@
However, ControlNet trained with dilated masks implicitly assumes the presence of dilation in the provided mask, even for precise masks. Fig. 3 reports the CR referring to precise mask $\boldsymbol{m}_{0}$ when ControlNet-$\boldsymbol{m}_{r}$ is conditioned on dilated masks $\boldsymbol{m}_{r}$. As ControlNet-$\boldsymbol{m}_{r}$ keeps a CR on $\boldsymbol{m}_{0}$ over $0.5$ until $r>20$, suggesting that models infer and follow precise contours based on the provided $\boldsymbol{m}_{r}$. Examples are provided in the Appendix. This effect diminishes as the radius $r$ increases, and ControlNet-$\boldsymbol{m}_{r}(r\geq 40)$ shows weak contour-following ability and prompts object positions via masks.
@@
Moreover, training ControlNet with severely deteriorated masks, e.g., bounding-box masks, yields poorer CLIP-score and FID with precise masks as shown in Fig. 2. Examples in Fig. 7 illustrate that ControlNet-$\boldsymbol{m}_{\infty}$ probably misinterprets one whole object mask as multiple small objects. This reminds us that too weak contour instructions can also cause problematic images and deviate from user intentions. So, it is worthwhile to develop a method to make contour-following ability more controllable, thereby fulfilling robust interpretation with inexplicit masks of varying precision.
@@
Figure 3: Illustration of the inductive bias of ControlNet-$\boldsymbol{m}_{r}$ conditioned on $\boldsymbol{m}_{r}$, where high CR on $\boldsymbol{m}_{0}$ indicates models implicitly learn the dilation radius $r$.
4.3 Impact of Hyperparameters
@@
Figure 4: Metric curves of (a) CFG scale, (b) conditioning scale, and (c) condition injection strategy. Red denotes the performance on the precise mask $\boldsymbol{m}_{0}$, and blue denotes the bounding-box mask $\boldsymbol{m}_{\infty}$
@@
As a common practice, adjusting hyperparameters like CFG scales is useful to achieve satisfactory generation results. Therefore, we investigate how hyperparameters impact the contour-following ability and reveal several guidelines.
@@
Specifically, we test the vanilla ControlNet (ControlNet-$\boldsymbol{m}_{0}$), under two extreme condition cases, i.e., precise mask $\boldsymbol{m}_{0}$ and bounding-box mask $\boldsymbol{m}_{\infty}$. Three core hyperparameters are considered, i.e., CFG scale $\lambda$, conditioning scale $\omega$, and condition injection strategy. Quantitative results are revealed in Fig. 4. Examples are presented in the Appendix.
@@
Under both conditions, ControlNet maintains high CR scores under all CFG scales, showing little impact on contour-following ability. This is because conditioning signals are added to both $\boldsymbol{\epsilon}_{\theta}(\cdot)$ and $\boldsymbol{\epsilon}_{\theta}(\cdot,\boldsymbol{c})$ in Eq. 2. Moreover, performance on precise mask $\boldsymbol{m}_{0}$ always surpasses that on $\boldsymbol{m}_{\infty}$, indicating precise control inputs improve image quality. The trends of CLIP-Score and FID curves align with common sense, where a relatively higher $\omega$ brings better fidelity and text faithfulness.
@@
The conditioning scale decides the strength of control signals as Eq.1. A higher $\lambda$ imposes stronger instructions, consistent with increasing CR scores under both conditions. However, the behavior diverges on CLIP-Score and FID when increasing $\lambda$. While both CLIP-Score and FID almost consistently increase on $\boldsymbol{m}_{0}$, its performance on $\boldsymbol{m}_{\infty}$ increases slightly at the very beginning but then declines dramatically, showing a large performance gap with $\boldsymbol{m}_{0}$. This reveals a trade-off between image quality and spatial control with bounding-box masks, where $\lambda\in[0.5,0.7]$ are recommended. Finding suitable $\lambda$ for masks of varying precision usually requires tens of attempts.
@@
We divide the sampling steps into ten stages during the reverse sampling phase for experiments. As shown in Fig. 4(c), all stages contribute to contour instructions during the generation process under both conditions. The first $10\%\sim 40\%$ steps hold the key, consistent with observations in Mou et al. (2023). In addition, a similar divergence of CLIP-Score and FID curves is observed between the two conditions, further declaring the performance gap between explicit and inexplicit masks.
@@
In summary, while the CFG scale has little impact on the contour-following ability, smaller conditioning scales and discarding conditions at early reverse sampling stages help to relieve contour instructions. This strategy helps to improve image fidelity and text faithfulness for inexplicit masks. However, subtle changes to such hyperparameters may lead to dramatic changes in image appearance (see Appendix). Hence, it is tricky to search for the ideal combination of hyperparameters for each image, and it takes risks of violating spatial control for a trade-off.
@@
Figure 5: Illustrations on the difference between explicit and inexplicit masks for control. The contour-following ability induces strong priors from contours, causing conflicts with textual prompts.
4.4 Empirical Analysis
@@
To clarify the difference between explicit and inexplicit mask controls, we conduct comparisons between ControlNet-$\boldsymbol{m}_{0}$ and ControlNet-$\boldsymbol{m}_{\infty}$ on precise masks $\boldsymbol{m}_{0}$ and bounding-box masks $\boldsymbol{m}_{\infty}$ in Fig. 5. By removing textual prompts, we find that the strong contour-following ability of vanilla ControlNet derives intense shape priors from contours, thus it infers specific objects from the mask. As inexplicit masks introduce incorrect shape priors, they result in conflicts with textual prompts or violation of spatial control. For example, bounding-box masks are misinterpreted as the door in Fig. 5. This is why vanilla ControlNet fails to interpret inaccurate contours with correct content. VisorGPT Xie et al. (2023) also validates that conditional generative models implicitly learn visual priors from the data, such as size and shape. If spatial conditions deviate from that prior, it would lead to artifacts and incorrect contents. In contrast, ControlNet-$\boldsymbol{m}_{\infty}$ relies less on the shape priors from the conditional mask but prompts object locations for spatial control. This leads to improved text alignment and robustness on inexplicit masks. However, ControlNet-$\boldsymbol{m}_{\infty}$ exhibits weak control with precise masks. So, it is necessary to develop a method to adapt ControlNet to masks of varying precision.
5 Shape-aware ControlNet
@@
To enhance ControlNet’s ability to interpret inexplicit masks robustly instead of blindly adhering to contours, we introduce a novel deterioration estimator and a shape-prior modulation block, namely Shape-aware ControlNet, as depicted in Fig. 6.
Figure 6: The overall architecture of Shape-aware ControlNet. It contains 1) a deterioration estimator to assess the deterioration ratio of inexplicit masks, and 2) a shape-prior modulation block to modulate this ratio to ControlNet to adjust the contour-following ability for robust spatial control with inexplicit masks.
5.1 Deterioration Estimator
@@
We first introduce the deterioration ratio $\rho$, which depicts the gap between the inexplicit masks and explicit shape priors of specific objects, formulated as,
$\rho=\frac{|S(\boldsymbol{m}_{r})-S(\boldsymbol{m}_{0})|}{|S(\boldsymbol{m}_{\infty})-S(\boldsymbol{m}_{0})|},$ (5)
@@
5.2 Shape-prior Modulation Block
Taking the shape prior $\rho$ as an extra condition, we design a shape-prior modulation block inspired by Karras et al. (2019). Specifically, we encode $\rho$ into Fourier embedding. Then, we employ a hypernetwork Ha et al. (2016) to modulate shape priors to zero convolution layers in ControlNet as illustrated in Fig. 6(c). Supposing a zero convolution $Z(\cdot;w)$ parameterized by $w$ and ControlNet encoder feature $F_{ij}$, the modulated feature $F^{\prime}_{ij}$ is computed as,
$F^{\prime}_{ij}=Z(F_{ij};(1+\Delta W)\cdot w),$ (6)
where $\Delta W=H(\rho)$, and $H(\cdot)$ is a hypernetwork constructed by multiple linear layers.
6 Experiments
Figure 7: Visualized comparison between the vanilla ControlNet and Shape-aware ControlNet with masks at varying levels of precision.
6.1 Implementation Details
The baselines including ControlNet-$\boldsymbol{m}_{0}$ and ControlNet-$\boldsymbol{m}_{\infty}$ are implemented as § 4.1. We inherit the same setting to train our shape-prior modulation block on COCO-LVIS for 10 epochs from scratch with a learning rate of $51e-$ 5, where the dilation radius is uniformly sampled from 00 to $80$. The deterioration estimator is trained for another 10 epochs individually with gradient-detached features from ControlNet.

### v3

@@ 4.2. Impact of Inexplicit Masks
To clarify concerns about ControlNet’s performance on inexplicit masks, we first imitate inaccurate control signals by dilating object masks $\boldsymbol{m}$ with progressively increasing radius $r(r\geq 0)$. The dilated mask is denoted as $\boldsymbol{m}_{r}$, where $\boldsymbol{m}_{0}$ represents masks with precise contours. Considering that bounding-box masks only convey object position and size without identifiable contours, we take an intersection over the dilated mask $\boldsymbol{m}_{r}$ and the bounding-box mask denoted as $\boldsymbol{b}$, to construct inexplicit masks of various degrees, formulated as,
(4) $\boldsymbol{m}_{r}=Dilate(\boldsymbol{m},r)\cap\boldsymbol{b},$
@@
where the bounding-box mask $\boldsymbol{b}$ can also be expressed as the extreme case $\boldsymbol{m}_{\infty}$ ($\boldsymbol{b}\subset\boldsymbol{m}_{\infty}$). In the rest of the paper, we use $\boldsymbol{m}_{\infty}$ to denote the bounding-box mask. As $r$ grows large, the masks lose detailed shape information gradually.
@@
We use the notation ControlNet-$\boldsymbol{m_{r}}$ to denote ControlNet trained with deteriorated masks $\boldsymbol{m}_{r}$, where the vanilla ControlNet trained with precise masks is ControlNet-$\boldsymbol{m_{0}}$. We test these models on masks of varying deterioration degrees to explore the performance of ControlNet with inexplicit masks and its contour-following ability. Fig. 2 presents the experimental results. Here are two main observations.
@@
The vanilla ControlNet, i.e., ControlNet-$\boldsymbol{m}_{0}$, strictly adheres to the provided outlines, where inexplicit masks would cause severe performance degradation. As shown in Fig. 1, the generated images are faithful to the contours of the control inputs under various control modalities, which is the key to excellent spatial controllable generation. One underlying reason is supposed to be that edges provide a shortcut to reconstruct the input images. Therefore, it exhibits strong contour instructions during the controllable text-to-image generation process of ControlNet.
@@
However, we notice its contour-following ability is blind, which takes no account for inaccurate contours and maintains a high average CR of over 60% in Fig. 2. While it allows minor editing by manipulating masks, this property imposes severe problems on image realism when the contours are inaccurate (e.g. user scribbles). Similar observations are also noted in (Voynov et al., 2023; Bhat et al., 2023; Singh et al., 2023), but an in-depth analysis of this fact is neglected. For example, ControlNet tends to interpret rectangle masks as boards rather than animals specified in the prompt, causing incorrect spatial layouts and distorted objects. This fact causes a degradation of 1.35 CLIP-score and 8.62 FID in Fig. 2. Thus, precise masks with accurate contours are necessary for high-quality image generation. However, obtaining masks with accurate contours is challenging, especially for non-expert users.
@@
Training ControlNet with deteriorated masks also converges and exhibits high robustness on masks of varying deterioration degrees with a sacrifice of contour instructions. We further train a series of ControlNet, i.e. ControlNet-$\boldsymbol{m}_{r}$, on progressively dilated masks $\boldsymbol{m}_{r}$. Surprisingly, ControlNet converges with all conditions, even on the extreme bounding-box masks. As depicted in Fig. 2, training with deteriorated masks consistently improves ControlNet’s robustness in interpreting inexplicit masks with better CLIP-score and FID at the cost of weak contour-following ability.
@@
However, ControlNet trained with dilated masks implicitly assumes the presence of dilation in the provided mask, even for precise masks. Fig. 3 reports the CR referring to precise mask $\boldsymbol{m}_{0}$ when ControlNet-$\boldsymbol{m}_{r}$ is conditioned on dilated masks $\boldsymbol{m}_{r}$. ControlNet-$\boldsymbol{m}_{r}$ keeps a CR on $\boldsymbol{m}_{0}$ over $0.5$ until $r>20$, suggesting that models infer and follow precise contours based on $\boldsymbol{m}_{r}$. Visualized examples are provided in Appendix Fig. S7. This inductive bias diminishes as the radius $r$ increases, and ControlNet-$\boldsymbol{m}_{r}(r\geq 40)$ shows weak contour-following ability and prompts object positions via masks.
@@
Moreover, training ControlNet with severely deteriorated masks, e.g., bounding-box masks, yields poorer CLIP-score and FID with precise masks as shown in Fig. 2. Examples in Fig. 7 illustrate that ControlNet-$\boldsymbol{m}_{\infty}$ probably misinterprets one whole object mask into multiple small objects. This reminds us that too weak contour instructions can also cause problematic images and deviate from user intentions. So, it is worthwhile to develop a method to make contour-following ability more controllable, thereby fulfilling robust interpretation with inexplicit masks of varying degrees.
@@
Figure 3. Illustration of the inductive bias of ControlNet-$\boldsymbol{m}_{r}$ conditioned on $\boldsymbol{m}_{r}$, where high CR on $\boldsymbol{m}_{0}$ indicates models implicitly learn the dilation radius $r$.
4.3. Impact of Hyperparameters
@@
Figure 4. Metric curves of (a) CFG scale, (b) conditioning scale, and (c) condition injection strategy for the vanilla ControlNet. Red denotes the performance on the precise mask $\boldsymbol{m}_{0}$, and blue denotes the bounding-box mask $\boldsymbol{m}_{\infty}$.
Figure 5. The overall architecture of Shape-aware ControlNet. It contains 1) a deterioration estimator to assess the deterioration ratio of inexplicit masks, and 2) a shape-prior modulation block to modulate this ratio to ControlNet to adjust the contour-following ability for robust spatial control with inexplicit masks.
@@
As a common practice, adjusting hyperparameters like CFG scales is a useful tool to achieve satisfactory generation results in T2I applications. Therefore, we investigate how hyperparameters impact the contour-following ability, especially when the conditional inputs are inexplicit, and reveal several guidelines.
@@
Specifically, we test the vanilla ControlNet (i.e., ControlNet-$\boldsymbol{m}_{0}$), under two extreme condition cases, i.e., precise mask $\boldsymbol{m}_{0}$ and bounding-box mask $\boldsymbol{m}_{\infty}$. Three core hyperparameters are considered, i.e., CFG scale $\lambda$, conditioning scale $\omega$, and condition injection strategy. Quantitative results are revealed in Fig. 4. For page limitations, visualized examples are presented in Appendix Fig. S8.
@@
Under both conditions, ControlNet maintains high CR scores under all CFG-scale settings, showing little impact on contour-following ability. This is because conditioning signals are added to both $\boldsymbol{\epsilon}_{\theta}(\cdot)$ and $\boldsymbol{\epsilon}_{\theta}(\cdot,\boldsymbol{c})$ in Eq. 2. Moreover, performance on precise mask $\boldsymbol{m}_{0}$ always surpasses that on $\boldsymbol{m}_{\infty}$, indicating precise control inputs improve image quality. The trends of CLIP-Score and FID curves align with common sense, where a relatively higher $\omega$ brings better fidelity and text faithfulness.
@@
The conditioning scale decides the strength of control signals as Eq. 1. A higher $\lambda$ imposes stronger instructions, consistent with increasing CR scores under both conditions. However, the behavior diverges on CLIP-Score and FID curves when increasing $\lambda$. While both CLIP-Score and FID almost consistently increase on $\boldsymbol{m}_{0}$, its performance on $\boldsymbol{m}_{\infty}$ increases slightly at the very beginning but then declines dramatically, showing a large performance gap with $\boldsymbol{m}_{0}$. This reveals a trade-off between image quality and spatial control with bounding-box masks, where $\lambda\in[0.5,0.7]$ are recommended. Finding suitable $\lambda$ for masks of varying deterioration degrees usually requires tens of attempts.
@@
We divide the sampling steps into ten stages during the reverse sampling phase for experiments. As shown in Fig. 4(c), all stages contribute to contour instructions during the generation process in both cases. The first $10\%\sim 40\%$ steps hold the key, consistent with observations in T2I-Adapter (Mou et al., 2024). In addition, a similar divergence of CLIP-Score and FID curves is observed between the two conditions, further declaring the performance gap between explicit and inexplicit masks.
@@
In summary, while the CFG scale $\omega$ has little impact on the contour-following ability, a smaller conditioning scale $\lambda$ and discarding conditions at early reverse sampling stages help to relieve contour instructions. These strategies help to relieve the degradation of image fidelity and text faithfulness for inexplicit masks and achieve a trade-off between image quality and spatial control. However, subtle changes to such hyperparameters may lead to dramatic changes in image appearance (see Appendix Fig. S8). Hence, it is still tricky to search for the ideal combination of hyperparameters for each image, and it also takes risks of violating spatial control.
@@
Figure 6. Illustrations on the difference between explicit and inexplicit masks for control. The contour-following ability induces strong priors from contours, which causes conflicts with textual prompts and misinterprets conditional inputs.
4.4. Empirical Analysis
To clarify the difference between explicit and inexplicit mask controls, we compare ControlNet-$\boldsymbol{m}_{0}$ with ControlNet-$\boldsymbol{m}_{\infty}$ on precise masks $\boldsymbol{m}_{0}$ and bounding-box masks $\boldsymbol{m}_{\infty}$ in Fig. 6.
@@ 4.4. Empirical Analysis
By removing textual prompts, we find that the strong contour-following ability of vanilla ControlNet derives intense shape priors from contours, thus it tends to infer specific objects from the mask directly. Since inexplicit masks introduce incorrect shape priors, they will cause conflicts with textual prompts or violations of spatial control for generation. For example, bounding-box masks are misinterpreted as the door in Fig. 6. This is why the vanilla ControlNet fails to interpret inaccurate contours with the correct content. VisorGPT (Xie et al., 2023) also validates that conditional generative models implicitly learn visual priors from the data, such as size and shape. If spatial conditions deviate from that prior, it would lead to artifacts and incorrect contents. In contrast, ControlNet-$\boldsymbol{m}_{\infty}$ relies less on the shape priors from the conditional mask and prompts object locations for spatial control. This leads to improved text alignment and robustness on inexplicit masks. However, ControlNet-$\boldsymbol{m}_{\infty}$ exhibits weak control with precise masks. So, it is necessary to develop a method to adapt ControlNet to deteriorated masks of varying degrees.
5. Shape-aware ControlNet
@@
To improve ControlNet’s interpretation of inexplicit masks rather than blindly adhering to contours, a novel deterioration estimator and a shape-prior modulation block are proposed, namely Shape-aware ControlNet, as depicted in Fig. 5. Our main insight is to introduce the shape prior explicitly to the model for better controllability.
5.1. Deterioration Estimator
@@
We first introduce a deterioration ratio $\rho$ to depict the gap between inexplicit masks and explicit shape priors of objects, formulated as,
(5) $\rho=\frac{|S(\boldsymbol{m}_{r})-S(\boldsymbol{m}_{0})|}{|S(\boldsymbol{m}_{\infty})-S(\boldsymbol{m}_{0})|},$
@@
5.2. Shape-prior Modulation Block
Taking the shape prior $\rho$ as an extra condition, we design a shape-prior modulation block inspired by StyleGAN (Karras et al., 2019). Specifically, we encode $\rho$ into Fourier embedding. Then, we employ a hypernetwork (Ha et al., 2016) to modulate shape priors to zero convolution layers in ControlNet as illustrated in Fig. 5(c). Supposing a zero convolution $Z(\cdot;w)$ parameterized by $w$ and ControlNet encoder feature $F_{ij}$, the modulated feature $F^{\prime}_{ij}$ is computed as,
(6) $F^{\prime}_{ij}=Z(F_{ij};(1+\Delta W)\cdot w),$
where $\Delta W=H(\rho)$, and $H(\cdot)$ is a hypernetwork constructed by multiple linear layers with normalization.
6. Experiments
Figure 7. Visualized comparison between the vanilla ControlNet and our Shape-aware ControlNet with masks at varying deterioration degrees. Our method not only follows explicit masks but also interprets inexplicit masks robustly.
6.1. Implementation Details
The baselines including ControlNet-$\boldsymbol{m}_{0}$ and ControlNet-$\boldsymbol{m}_{\infty}$ are implemented as § 4.1. We inherit the same setting to train our shape-prior modulation block on COCO-LVIS dataset for 10 epochs from scratch with a learning rate of $51e-$ 5. The dilation radius $r$ is uniformly sampled from 00 to $80$. The deterioration estimator is trained for another 10 epochs individually with gradient-detached features from the ControlNet encoder.